OpenStudy (anonymous):

Robert borrowed $750 from his parents to make some repairs on his car. He promised to repay the loan by giving his parents at least $67 from his paycheck each week. Assuming Robert sticks to his promise, which of the following is not a possible scenario for the amount of debt Robert has remaining?

OpenStudy (anonymous):

following...

OpenStudy (anonymous):

After 5 weeks, Robert is $150 in debt. After 6 weeks, Robert is $200 in debt. After 7 weeks, Robert is $250 in debt. After 8 weeks, Robert is $300 in debt

OpenStudy (callisto):

Consider the minimum payment Week interest paid loan (at most) (x = total loan - interest paid) 1 67 750 - 67 = 683 2 67 x 2 = 134 750 - 134 = 616 3 67 x 3 = 201 750 - 201 = 549 4 67 x 4 = 268 750 - 268 = 482 5 67 x 5 = 335 750 - 335 = 415 6 67 x 6 = 402 750 - 402 = 345 7 67 x 7 = 469 750 - 469 = 281 8 67 x 8 = 536 750 - 536 = 214 The amount of the loan can be less than x as he may pay more than 67 in any weeks. However, it cannot be greater than x as he keeps his promise to pay the minimum amount of 67. Can you get the answer now?

OpenStudy (anonymous):

Answer above is perfect.

OpenStudy (callisto):

Thanks!!!1

OpenStudy (anonymous):

IT SEEMS LIKE 8 WEEKS IS THE ANSWER.